OpenStudy (anonymous):

what is the probability of either a boy or girl being born when the probability of a boy being born equals .50 or 1/2 as does the probability of a girl being born?

OpenStudy (anonymous):

50/50

OpenStudy (anonymous):

they each have a 50% chance

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

simple. the probability of a boy being born equals .50 or 1/2 as does the probability of a girl being born, so they each have a 50% chance

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

what about for 3 boys?

OpenStudy (anonymous):

what do you mean for 3 boys, 3 boys out of what?

OpenStudy (anonymous):

for a randomly selected family with 3 children what's the probability of 3 boys ?

OpenStudy (anonymous):

50%, if we are using the same probability as before. if there is a 50% chance of a boy and a 50% chance of a girl, then te probability of three boys is 50%

OpenStudy (anonymous):

same as if it where 3 girls, 50%

OpenStudy (anonymous):

you get it?

OpenStudy (zarkon):

the probability of 3 boys is not 50%

OpenStudy (anonymous):

how? I thought it would be 1/3 because the family wants 3 children, and all 3 to be boys?

OpenStudy (zarkon):

the probability that one has a boy is 1/2 (for each boy). if the births are independent then the probability of 3 consecutive boys is (1/2)(1/2)(1/2)=1/8

OpenStudy (anonymous):

same for 3 girls right? and how about either a boy or a girl?

OpenStudy (zarkon):

BBB BBG BGB GBB BGG GBG GGB GGG the 8 above are the only possibilities...1/8 are BBB and 1/8 are GGG

OpenStudy (anonymous):

what about either a boy or a girl? What should the equation setup look like?

OpenStudy (zarkon):

P(all boys)=1/8 P(all girls)=1/8 P(all boys or all girls)=P(all boys)+P(all girls)=1/8+1/8=2/8=1/4

OpenStudy (anonymous):

what about neither?

OpenStudy (anonymous):

?